using formula w=0.00185h^2.67 where w is a boy's weight in pounds and h is his height in inches, estimate the height of a boy whose weight is 100 pounds.
Answer by checkley77(12844)

w=0.00185h^2.67
100=0.00185H^2.67
H^2.67=100/0.00185
H^2.67=54,054
2.67\/54,054=59.24 INCHES TALL.